OpenStudy (anonymous):

(porportions) 4 2 - = - 3x 3

OpenStudy (aravindg):

Take LCM.

OpenStudy (anonymous):

huh

OpenStudy (aravindg):

a/b=c/d ad=bc

OpenStudy (anonymous):

2?

OpenStudy (aravindg):

Yes Good work!

OpenStudy (anonymous):

thanks now i get it

OpenStudy (aravindg):

yw :) Click best response for answers that help. Thank you.